OpenStudy (anonymous):

help? find the values of k which will make the roots of the equation real: kb^2-9b+4=0

hartnn (hartnn):

h @Noodles24 \(\Huge \mathcal{\text{Welcome To OpenStudy}\ddot\smile} \) compare your quadratic equation with \(Ab^2+Bb+C=0\) the roots will be equal when \(B^2-4AC=0\)

hartnn (hartnn):

so find A,B,C first

OpenStudy (anonymous):

A=k, B=9, C=4?

hartnn (hartnn):

B = -9 isn't it ?

OpenStudy (anonymous):

yes you are right, my bad

hartnn (hartnn):

because the 2nd term is -9b ok, so now calculate B^2-4AC ?

hartnn (hartnn):

\(81 - 4\times 4\times k =0\) can you find k from here ?

OpenStudy (anonymous):

yes thank you very much, you were a ton of help! :)

hartnn (hartnn):

actually, for real roots, its \(B^2-4AC >0\)

hartnn (hartnn):

so, please solve \(81-16k \ge 0\) and not =

OpenStudy (anonymous):

oh ok

hartnn (hartnn):

***\(B^2-4AC \ge 0\)

hartnn (hartnn):

let me know what you get for 'k' so that i can verify :)

OpenStudy (anonymous):

k is less than or equal to 81/16?

hartnn (hartnn):

good! absolutely correct :)

OpenStudy (anonymous):

yay! thank you so much :)

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

could you help me with another problem?

OpenStudy (anonymous):

i am doing a regents review sheet and a couple are hard >_<

hartnn (hartnn):

sure :) ask.

OpenStudy (anonymous):

determine the constant that must be added to the given expression to complete the square and express the resulting trinomial as the square of a binomial.

OpenStudy (anonymous):